QUESTION 13 3.5 points Save Answer Solve the equation 4 A B E QUESTION 14 3.5 points Save Answer Find the slope of the line tSteps too so I can understand it please.
Answer #1

Given that 4 + 2 - 2 h 13 40+32e - 10 Jez > 4x+30-10=0 For Zeros of a, comparing wood butc20 with x = -b orsac - where a=4 b=3 Given that 2x4 9/4x? = (4x4) (4x97% 1/3 G4) 4 439) is (69) 4/13 % x 26 mtn 33 - 4 X 7 [.fabiacam.com anar=a 12 Camin= amn (